OpenStudy (anonymous):

Write an equation of the line the passes through (-5,8) and (-3,-8)

OpenStudy (mimi_x3):

Well, first you have to find the gradient in which the formula is: \[y = \frac{y_2-y_1}{x_2-x_1}\] Then the equation of the line is: \[y=mx+b\] This is where you sub the \(y\) value; \(m\) value, and \(x\) value that is given to find \(b\). After that the equation of the line passing through the points \(x_1,y_1),(x_2,y_2\) is found! Try it; if you haven't!

OpenStudy (mimi_x3):

Well, then tell me what you get. Or tell me where you are having problems if you have not solved it.

OpenStudy (anonymous):

I solved It But I Dont Know How To Work it At All @Mimi_x3

OpenStudy (mimi_x3):

Where are you having problems?

OpenStudy (anonymous):

All Of It @Mimi_x3

OpenStudy (mimi_x3):

\((-5, 8)\) \((-3,-8)\) => \((x_1,y_1),(x_2,y_2)\) Then sub it in; into the gradient formula given.

OpenStudy (anonymous):

i Know that part i just dont know to put it in a formula @Mimi_x3

OpenStudy (mimi_x3):

Can you show me what you have tried?

OpenStudy (anonymous):

I got 2/16 @Mimi_x3

OpenStudy (mimi_x3):

\[y = \frac{-8-8}{-3-5} = \frac{-16}{-8}\]

OpenStudy (mimi_x3):

That should be \(m\) not \(y\)

OpenStudy (anonymous):

now how do you put in the equation @Mimi_x3

OpenStudy (mimi_x3):

the equation \(y=mx+b\)

OpenStudy (mimi_x3):

you tell me what is the \(y\) value?

OpenStudy (mimi_x3):

from \(( (-5,8) and (-3,-8)\)